Find the volume of the solid generated by revolving about the x-axis by the curve, y2 = 9x and the line y = 3x


1
Expert's answer
2021-06-09T11:37:01-0400

"V=2x \int_2^1 (2-x)(x^3+x+1-8)dx"

"=2x\int^2_0(-x^4+2x^3-x^2+2x)dx"

"=[-\frac{x^5}{5}+\frac{x^4}{4}-\frac{x^3}{3}+x^2]_0^1"

"=2\pi[\frac{-1}{5}+\frac{1}{2}+\frac{1}{3}+1]"

"=\frac{29\pi}{15}"